# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ A.6 Implicitly Dealing with the Terminal Vertex
Recall that these arcs either represent donations that are made to the DDWL or donors becoming bridge donors in future matching rounds. The associated variables are not required when the associated weights $w_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$ are zero. Moreover, even when this is not the case, we can omit these variables by implicitly modelling $\tau$ while still accounting for the corresponding weights.
@@
  • For EEF-CHAIN-EXP and EEF-CHAIN-MTZ we can remove from each $\mathcal{A}^{s}$ all arcs in $\mathcal{A}_{\tau}$ and omit the associated variables. Moreover, the equality sign in constraints (48) should be replaced with $\leq$”, and the factor $L$ on the right-hand side of constraints (49) should be replaced by $L-1$.
@@
A.7 Dealing with an Unbounded Maximum Chain Length
@@
A.8 Results on Different Subsets of Instances
To complement the main computational results presented in Section 4.2, we present here the results of some final experiments in which we explore the impact of different instance parameters. We focus on the most effective methods as indicated in Section 4.2, namely our implementations of combined models CF-CYCLE+PIEF-CHAIN, HCF-CYCLE+PIEF-CHAIN and PIEF-CYCLE+PIEF-CHAIN, as well as third-party methods JL-BNP and JL-BNP-PICEF. Furthermore, in the unweighted case, we consider the RCVF implementations of the combined models, whereas we consider the standard implementations in the weighted case.
@@ A.8 Results on Different Subsets of Instances
Apart from whether an instance is weighted or not, the key parameter that affects the hardness of an instance is the number of RDPs $|\mathcal{R}|$. Whereas all methods manage to solve all unweighted and most weighted instances up to size $200$, each of the methods starts to have difficulties for instances of order $500$ and higher, with considerable room for improvement for weighted instances with ||$|\mathcal{R}|\in\{750,1000\}$.
@@
Finally, the performance of most methods becomes worse as the cycle length limit $K$ and the chain length limit $L$ increase. As could be expected from the worst-case upper bounds on the number of variables and constraints presented in Table 8 in Appendix A.4, the performance of cycle model CF-CYCLE depends on $K$ most strongly, while the dependence on $K$ is lower for HCF-CYCLE and even less for PIEF-CYCLE. As a result, PIEF-CYCLE+PIEF-CHAIN becomes relatively more dominant with respect to CF-CYCLE+PIEF-CHAIN and HCF-CYCLE+PIEF-CHAIN as $K$ increases. Interestingly, the performance of JL-BNP and JL-BNP-PICEF does not strongly depend on $K$ in the unweighted case, which aligns with the fact that these methods rely on column generation. However, in the weighted case, the performance of JL-BNP and JL-BNP-PICEF does notably drop when $K$ increases from $3$ to $4$.

### v2

@@ Appendix G Implicitly Dealing with the Terminal Vertex
Recall that these arcs either represent donations that are made to the DDWL or donors becoming bridge donors in future matching runs. The associated variables are not required when the associated weights $w_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$ are zero. Moreover, even when this is not the case, we can omit these variables by implicitly modelling $\tau$ while still accounting for the corresponding weights.
@@
  • For EEF-CHAIN-EXP and EEF-CHAIN-MTZ we can remove from each $\mathcal{A}^{s}$ all arcs in $\mathcal{A}_{\tau}$ and omit the associated variables. Moreover, the equality sign in constraints (48) should be replaced with $\leq$”, and the factor $L$ on the right-hand side of constraints (49) should be replaced by L1L-1.
@@
Appendix H Dealing with an Unbounded Maximum Chain Length
@@
Appendix I Results of RCVF Implementations
To complement the main computational results presented in Section 4.2, we present in Table 9 the performance of the most relevant model combinations when solved using RCVF. Specifically, we report the number of instances solved to optimality within the time limit (column “#opt”) and the average CPU time across all instances (column “t”), in both the unweighted and the weighted case. To provide a comparison, the numbers in brackets indicate the performance of the standard implementations of these models. In addition, we present for each model the average number of iterations required by RCVF (measured as the number of calls to the ILP solver, column “#iter”) and the percentage of variables remaining in the reduced ILP model at the final RCVF iteration relative to the full model (column “%vrem”). To compute “#iter” and “%vrem” we only considered the 2298 instances where all models could be solved to optimality for both the unweighted and weighted variant of the instance.
Table 9: Performance of the RCVF implementations across 2160 unweighted and 2160 weighted instances.
unweighted instances weighted instances
model #opt t #iter %vrem #opt t #iter %vrem
Note that the number of RCVF iterations is mostly consistent across all models for a given instance, with the PIEF-CHAIN models sometimes requiring a few more iterations compared to the CF-CHAIN models. This is a direct result of the fact that the number of RCVF iterations required is determined solely by the absolute gap between the optimal value of an ILP model and its LP relaxation, and these gaps are identical or almost identical for the models considered here (see also Section 3.8 and Table 5). This also explains why weighted instances require more RCVF iterations than unweighted ones, as the absolute gap is larger for weighted instances compared to unweighted ones. On the other hand, the reduction in the number of variables due to RCVF is substantially larger for the weighted case compared to the unweighted case. The reduction is also larger for the CF-CHAIN models compared to the PIEF-CHAIN models, in particular for weighted instances, explaining why the CF-CHAIN models still benefit from RCVF in that case. Finally, while we only report statistics on the number of iterations and percentage of remaining variables for solved instances, we note that the unweighted instances where the time limit is reached are characterised by a percentage of remaining variables that is typically larger than that of solved unweighted instances. Conversely, the weighted instances where the time limit is reached are characterised by a number of iterations that is generally larger compared to that of solved weighted instances.
Appendix J Results on Different Subsets of Instances
Finally, we present here the results of some last experiments in which we explore the impact of different instance parameters. We focus on the most effective methods as indicated in Section 4.2 and Appendix I, namely our implementations of combined models CF-CYCLE+PIEF-CHAIN, HCF-CYCLE+PIEF-CHAIN and PIEF-CYCLE+PIEF-CHAIN, as well as third-party methods JL-BNP and JL-BNP-PICEF. Furthermore, in the unweighted case, we consider the RCVF implementations of the combined models, whereas we consider the standard implementations in the weighted case.
@@ Appendix I Results of RCVF Implementations
Apart from whether an instance is weighted or not, the key parameter that affects the hardness of an instance is the number of RDPs $|\mathcal{R}|$. Whereas all methods manage to solve all unweighted and most weighted instances up to size $200$, each of the methods starts to have difficulties for instances of size $500$ and higher, with considerable room for improvement for weighted instances with $|\mathcal{R}|\in\{750,1000\}$.
@@
Finally, the performance of most methods becomes worse as the cycle length limit $K$ and the chain length limit $L$ increase. As could be expected from the worst-case upper bounds on the number of variables and constraints presented in Table 7 in Appendix E, the performance of cycle model CF-CYCLE depends on $K$ most strongly, while the dependence on $K$ is lower for HCF-CYCLE and even less for PIEF-CYCLE. As a result, PIEF-CYCLE+PIEF-CHAIN becomes relatively more dominant with respect to CF-CYCLE+PIEF-CHAIN and HCF-CYCLE+PIEF-CHAIN as $K$ increases. Interestingly, the performance of JL-BNP and JL-BNP-PICEF does not strongly depend on $K$ in the unweighted case, which aligns with the fact that these methods rely on column generation. However, in the weighted case, the performance of JL-BNP and JL-BNP-PICEF does notably drop when $K$ increases from $3$ to $4$.